Sanna rolls a ball up to another person on a smooth ramp 19.6 m above her. The
ball reaches the other person’s hands when it is travelling 4.9 m/s uphill. If the
ramp angle slows the ball down by 3.7 m/s each second it travels up the ramp, find
the initial velocity of the ball.
1
Expert's answer
2020-07-09T10:11:18-0400

According to the kinematic formula


"h=\frac{v_f^2-v_i^2}{2a}"


We have


"2ah=v_f^2-v_i^2\to v_i^2=v_f^2-2ah"


Each second the speed is decreased by "3.7 m\/s" . Acceleration is the change of velocity with time. So,


"a=-3.7\/1=-3.7m\/s^2"


"v_i^2=v_f^2-2ah\to v_i=\sqrt{v_f^2-2ah}=\sqrt{4.9^2-2\cdot (-3.7)\cdot 19.6}=13m/s"